OpenStudy (anonymous):

use the distributive property to solve the equation -3(m-6)=4

OpenStudy (anonymous):



OpenStudy (anonymous):

DISTRIBUTE!!!! 

OpenStudy (anonymous):

-3m + 18 = 4

OpenStudy (anonymous):

Got it yet? :D

OpenStudy (anonymous):

ya i think

OpenStudy (anonymous):

Great! :D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

how would i get m by itself?

OpenStudy (anonymous):

So... add -18 to both sides of the equation... -3m + 18 - 18 = 4 - 18

OpenStudy (anonymous):

-3m = -14

OpenStudy (anonymous):

Now divide both sides by -3: \[\large m = \frac{-14}{-3}\]

OpenStudy (anonymous):

so would m = -1

OpenStudy (anonymous):

No... not unless you're implying fourteen divided by three is 1 :D

OpenStudy (anonymous):

I think it's best to leave it as a fraction, 14/3

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

:)